Problem #3: Suppose the U.S. Treasury offers to sell you a bond for $747.25. No payments will be made until the bond matures 5 years from now, at which time it will be redeemed for S1,000. What interest rate would you earn if you bought this bond at the offer price?

How you would input this into a financial calculator?
Answer #1

PV = 747.25
FV = 1000
N= 5
PMT = 0
I/Y = (1000/747.25)^0.20 = 1.060002

SO the interest rate is 6.0002%.

In order to calculate in a financial calculator, provide below value

PV = 747.25
FV = 1000
N= 5
PMT = 0
hit CPT and then I/Y, it will give the interest rate as 6.0002%